OpenStudy (anonymous):

If sin(x)=-12/13 and sec(x)=13/5, then what is sin(2x)?

OpenStudy (anonymous):

being told that \[\sec(x)=\frac{13}{5}\] is the same as being told that \[\cos(x)=\frac{5}{13}\] now use \[\sin(2x)=2\sin(x)\cos(x)\] replacing \[\sin(x)\] with \[-\frac{12}{13}\]and \[\cos(x)\] with \[\frac{5}{13}\]